1) An alpha particle is whizzing by at a velocity determined by its KE of 5 Mev.

a ) What is the velocity of the alpha particle?

b) What is the magnetic field at the position of a nucleus 2nm away from the alpha particle at angle of 900 to the velocity of the alpha. Alpha particle has a charge of 2e but a mass of 4 protons.

2) A long solenoid of 80 turns and length 16 cm is known to have a field of 4x10 -3 T. What is the current in the coil?
Answer #1

Q.1 Part A) We know that

\\ K = \frac{1}{2}mv^2 \\ \\ v = \sqrt{\frac{2K}{m}} \\ \\ v = \sqrt{\frac{2 \times 5 \times 1.6 \times 10^{-13} \ \rm J}{4 \times 1.67 \times 10^{-28} \ \rm kg}} \\ \\ v = 4.89 \times 10^7 \ \rm m/s

Part B) Magnetic field at the center is given by

B = \frac{\mu_0 I}{2r}

Where

\\ I = \frac{q}{T} \\ \\ T = \frac{2 \pi r}{v} \\ \\ I = \frac{qv}{2\pi r}

B = \frac{ \mu_0 qv}{4\pi r^2}

B = \frac{ 4\pi \times 10^{-7} \ \rm T.mA^{-1} \times 2 \times 1.6 \times 10^{-19} \ \rm C \times 4.89 \times 10^{7}\ \rm m/s}{4\pi \times (2.0 \times10^{-9}\ \rm m)^2}

B = 0.39 \ \rm T

Q.2. We know that

\\ B = \mu_0 nI \\ \\ I = \frac{B}{\mu_0 n} \\ \\ I = \frac{BL}{\mu_0 N}

As n = N/L

\\ I = \frac{4 \times 10^{-3}\ \rm T \times 0.16 \ \rm m}{4 \pi \times 10^{-7}\ \rm T.mA^{-1} \times 80N} \\ \\ I = 6.37 \ \rm A